You are hired to build a geothermal power plant that absorbs heat from a hot water spring and discards heat into the surrounding air, which is at a temperature of 28°. The plant is designed for a maximum (Carnot) efficiency of 0.15.

What is the temperature, in degrees Celsius, of the hot water gushing from the spring?
Answer #1

Temperature of the hot water from the spring = Th

Temperature of the surrounding air = Tc = 28 oC = 28 + 273 K = 301 K

Carnot efficiency of the plant = \eta = 0.15

号-1=

301 0.15-1-

* = 0.85

Th = 354.12 K

Converting to degrees Celsius,

Th = 354.12 - 273

Th = 81.12 oC

Temperature of the hot water gushing from the spring = 81.12 oC